OpenStudy (anonymous):

HELP PLEASE: Given f(x) = 5x - 6 and g(x) = x2 - 5x + 6, find the function (f - g)(x). Use the ^ key for the exponent. Enter your answer as the following example: -2x^2+3x+10

OpenStudy (anonymous):

(f-g)(x)=(5x-6)-(x^2-5x+6) i think

OpenStudy (noelgreco):

(f-g)(x)=f(x)-g(x) Telling a person helping you how to format his answer is a little over the edge.

OpenStudy (anonymous):

@noelgreco that was how the question was..thank anyway @hheston thanks for your help!(: